OpenStudy (anonymous):

Rewrite with only sin x and cos x. cos 2x - sin x

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (ybarrap):

Is this $$ \cos^2x-\sin x\\ or\\ \cos 2x-\sin x $$ ?

OpenStudy (anonymous):

the second

OpenStudy (anonymous):

of which you provided

OpenStudy (ybarrap):

So, $$ \cos 2x=\cos^2x-\sin^2x $$ Combine this with \(\sin x\)

OpenStudy (anonymous):

not sure how thats why im here

OpenStudy (ybarrap):

$$ \cos 2x-\sin x\\ =\cos^2x-\sin^2x-\sin x $$ Make sense?

OpenStudy (anonymous):

yep simply multiplied each side by sin x

OpenStudy (ybarrap):

No. I found \(\cos 2x\) in terms of \(\cos x\) and \(\sin x\) and then added to \(\sin x\), which is already in the form we desire.

OpenStudy (anonymous):

alright, so the next step?